Define the operation $ as a$b= (a)(b) - a. What is the value of (3$5)$(5$3)?
Answer by Fombitz(32388)

a$b=a*b-a=a(b-1)
So,
3$5=3(5-1)=3(4)=12
5$3=5(3-1)=5(2)=10
(3$5)$(5$3)=12$10=12(10-1)=12(9)=108